OpenStudy (anonymous):

.1. Which property is demonstrated? 7 + 9 + 3 = 7 + 3 + 9 associative property of addition associative property of multiplication commutative property of addition commutative property of multiplication Question 2.2. Simplify 10 • 4 • 4 • 50 800 2000 4000 8000 Question 3.3. Which property is demonstrated? 72 + (28 + 93) = (72 + 28) + 93 associative property of addition associative property of multiplication commutative property of addition commutative

OpenStudy (anonymous):

Question1: Commutative property of Addition a+b =b+a In this case:9+3 = 3+9 Question2.2 8000 Question 3.3 Associative property of addition (a+b)+c = a+(b+c)